the cost of enclosing a rectangular garden with a fence all around ,at the rate of 75 paise per metre , is rs 300.if the length of the garden is 120m ,find the area of field in square metres.
Answer by richwmiller(17219)

I believe you have 100 paise to one rupee.
So 300 rupee has 30000 paise
total cost divided by cost per metre
30000/75=400 metres for the perimeter
one length =120 meteres
We have 2 lengths=240
400-240=160 =two widths
160 is two widths
160/2=80 width
so now we have a length of 120 an a width of 80
120*80 =area of the garden
I leave that for you to multiply